You inherit $13000 and you decide to invest the money in two different ways : one paying 10% and the other paying 14%. A year later your investments are worth %14580. How much did you originally invest in each account ?
Answer by nerdybill(7384)

You inherit $13000 and you decide to invest the money in two different ways : one paying 10% and the other paying 14%. A year later your investments are worth %14580. How much did you originally invest in each account ?
.
Let x = amount ($) invested at 14%
then
13000-x = amount($) invested at 10%
.
.14x + .10(13000-x) = 14580-13000
.14x + 1300-.10x = 1580
.04x + 1300 = 1580
.04x = 280
x = $7000 (amount invested at 14%)
.
amount invested at 10%:
13000-x = 13000-7000 = $6000